OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $1.64, and a college-ruled notebook for $1.17. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $66.02. How many of each type were sold?

OpenStudy (anonymous):

amount of graph paper sold = X

OpenStudy (anonymous):

amount of college-ruled sold = Y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the total cost of graph paper notebook sold is going to be 1.64 * X the total cost of college ruled notebook sold is going to be 1.17 * Y

OpenStudy (anonymous):

X + Y = 50 (total sold)

OpenStudy (anonymous):

1.64 * X + 1.17 * Y = 66.02

OpenStudy (anonymous):

solve for X and Y, and you're good to go.

OpenStudy (anonymous):

this is how i set it up too

OpenStudy (anonymous):

oh great! do you need help with how to solve it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets mutliply the first equation (x+y = 50) by 1.17

OpenStudy (anonymous):

you get 1.17x + 1.17y = 58.5

OpenStudy (anonymous):

subtract the second equation (1.68x + 1.17y = 66.02) by the new equation

OpenStudy (anonymous):

oh, i mean 1.64x + 1.17y = 66.02 no wonder it wasn't working when i was doing it by myself

OpenStudy (anonymous):

so subtract the new equation (1.17x + 1.17y = 58,5) from that equation

OpenStudy (anonymous):

ok i was wondering were that number came from

OpenStudy (anonymous):

ok so you get 0.47x = 7.52

OpenStudy (anonymous):

divide both sides by 0.47, and you get x = 16

OpenStudy (anonymous):

plug this into the first equation of x + y = 50, and you get y = 34

OpenStudy (anonymous):

so you sold 16graph paper notebooks and 34 college ruled notebooks.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem.:)